OpenStudy (anonymous):

Which mathematical sentence represents the solution for d in the equation 6e = ef + 3d?

OpenStudy (freethinker):

which? where are the choices?

OpenStudy (anonymous):

i know the answer is d=6e-ef all over 3 but i dont know how to figure it out

OpenStudy (freethinker):

huh? you're just putting the equation into words, why are you trying to solve it?

OpenStudy (anonymous):

because if i saw another equation like this, i wouldn't know how to get the answer. my teacher gave me this answer without explaining it.

OpenStudy (freethinker):

but you're trying to solve it in a manner that is not being asked you to do. if you are looking for d in 6e = ef + 3d? then you do the following: subtract ef on both sides of the equation (right-hand side and left-hand side) 6e - ef = 3d then you divide both sides by 3 (6e - ef)/3 = d

OpenStudy (anonymous):

thank you

OpenStudy (freethinker):

you can simplify it further by factoring: [ e (6 - f) ]/3 = d

OpenStudy (anonymous):

Thanks for your help! I may have a few more questions! :)